OpenStudy (anonymous):

Solve for x x^2 – 1 = x + 1

Parth (parthkohli):

Bring everything to one side. \(\Large \color{Black}{\Rightarrow x^2 - x - 2 = 0 }\) Quadratic equation!

OpenStudy (anonymous):

x^2 – 1 = x + 1 x^2-x-2=0

OpenStudy (anonymous):

(x-2)(x+1)

OpenStudy (anonymous):

fine?

Parth (parthkohli):

Fontez nooo

OpenStudy (anonymous):

no what? :)

OpenStudy (anonymous):

explain me please...

Parth (parthkohli):

You shouldn't give the whole answer!

OpenStudy (anonymous):

why? ahhhh?!!?!? why not?? openstudy ->sharing or not?

OpenStudy (anonymous):

news rules and conditions?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

strange, very strange...but ok. delete my account. not accord this conditions...bye and thanks.

Parth (parthkohli):

Ok, bye!